# Manolis Mylonas CERTH-ITI Thessaloniki, Greece

### v1

@@ 3.1 Problem formulation
Let us assume an input video of $N$ frames and a user-provided script (a free-form text description) of $M$ sentences outlining the content of the desired video summary; different sentences of the script may refer to different parts of the full-length video with varying visual and semantic content. Then, script-driven video summarization aims to select a subset of frames or fragments of the video that are both relevant to one or more sentences of the user-provided script, and important for an overall understanding of the full-length video. The selected frames/fragments should form a video summary that does not exceed a predefined length $L$. Following the paradigm in the video summarization literature [1], $L$ is typically set to $15\%$ of the full-length video duration.
@@
Another task that aims to spot parts of the video that are relevant to a freeform query is the task of temporal sentence grounding in videos [28] (a.k.a. video grounding). Nevertheless, this task tries to localize one target segment from the full-length video, that corresponds to a given sentence query. On the contrary, script-driven video summarization aims to generate a concise summary, and in this process it needs to select multiple video fragments that relate to the complex query, as well as other less relevant parts that are necessary for providing a synopsis of the full-length video.
3.2 Script-driven video summarization dataset
@@
To overcome observed weaknesses of small video summarization datasets, related to the sensitivity of the evaluation outcomes to different data splits (as shown in [29]) and the inability to select a well-trained model in the absence of a validation set (discussed in [1]), we focused on the VideoXum large-scale dataset for cross-modal video summarization [9]. VideoXum includes $14,000$ open-domain videos up to $12.5$ min. long with diverse visual content, from the ActivityNet Captions dataset [16]. Each video is associated with multiple ground-truth video summaries - in the form of frame-level binary scores which denote the inclusion (label “1”) or not (label “0”) of a frame in the video summary - that were obtained by $10$ different human annotators. The dense video captions from the ActivityNet Captions dataset that are considered as text summaries in [9] provide a high-level description of the full-length video and thus are more appropriate for multimodal generic video summarization purposes. The existence of multiple ground-truth summaries per video ($10$ in total) makes the VideoXum dataset well-suited for the script-driven video summarization task, as it allows to train a summarization method to generate different summaries for a given video, driven by the provided script about the content of each summary. Furthermore, its significantly large size allows to split the available videos into training, validation and test set. Such a data split supports the extensive training of a model using a very wide set of training samples, the selection of a well-trained model that exhibits good generalization capacity through the use of an adequately big validation set, and the robust evaluation of its performance using a large number of testing samples.
@@
To make VideoXum suitable for training and evaluation of script-driven video summarization methods, we extended it by producing natural language descriptions of the different ground-truth summaries that are available per video. For this, we employed the publicly-available [^1] https://huggingface.co/llava-hf/LLaVA-NeXT-Video-7B-hf state-of-the-art Large Multimodal Model LLaVA-NeXT-7B [10]. To align with the adopted video representation strategy in [9], we processed each ground-truth video summary using one frame per second and obtained the natural language descriptions using the following prompt: “Describe the important scenes in the video”. The maximum number of generated tokens was set equal to $200$, and a $4$-bit quantization was applied to reduce computational cost. Moreover, to allow performance comparisons with multimodal approaches for generic video summarization (which are compatible with our method in terms of the used data modalities), we produced natural language descriptions of the full-length videos, as well. Once again, we sampled these videos keeping one frame per second and generated descriptions using the LLaVA-NEXT-7B model and the same prompt.
Figure 2: Overview of the SD-VSum network architecture. Given an input video and a user script outlining the content of the desired summary, SD-VSum generates a summary conditioned on the script. Visual and text embeddings are obtained using the corresponding encoders of a pretrained multimodal vision and language model, and fused using a trainable cross-modal attention mechanism. The cross-modal embeddings in the output of this mechanism, that encode relationships between the different modalities and the position of the video frames, are forwarded to a trainable Transformer-based frame scoring mechanism, which computes frame-level importance scores. These scores are finally used by a frame/fragment selection component that forms the video summary based on a predefined temporal fragmentation of the video and a time-budget about the summary duration.
@@
Since some of the videos from the VideoXum dataset were not publicly-available during this work, the extended S-VideoXum dataset for query-driven video summarization differs from the original one in terms of number of videos, including data for $11,908$ videos. However, a statistical analysis on the obtained videos showed that our dataset maintains the same distribution of training, validation and test samples with the original VideoXum dataset. In particular, we use $6,782$ samples for training, $3,419$ for validation and $1,707$ for testing purposes. The generated natural language descriptions for the ground-truth summaries and the full-length videos of the S-VideoXum dataset, along with the extracted (CLIP-based) embeddings from visual and textual data and the used data splits in our experiments, are publicly-available on GitHub at: https://github.com/IDT-ITI/SD-Vsum.
4 Proposed Approach
4.1 Network architecture
@@
An overview of the SD-VSum network architecture is provided in Fig. 2. Assuming an input video of $N$ frames (following sampling), and a user script of $M$ sentences that outlines the content of the desired summary, SD-VSum produces a video summary conditioned on the script and a time-budget about its length. A set of visual embeddings ($\boldsymbol{X}=\{\boldsymbol{x_{n}}\}_{n=1}^{N}$) of size $D$ ($\boldsymbol{x_{n}}=\{x_{r}\}_{r=1}^{D}$) are obtained for the video frames, using the visual encoder of a pretrained multimodal embedding model. Similarly, a set of text embeddings ($\boldsymbol{Y}=\{\boldsymbol{y_{m}}\}_{m=1}^{M}$) of the same size ($\boldsymbol{y_{m}}=\{y_{r}\}_{r=1}^{D}$) are computed for the different sentences of the user script (one per sentence), using the text encoder of the aforementioned model. These embeddings are then given as input to a multi-head cross-modal attention mechanism, as shown in Fig. 3. Given the $h^{th}$ attention head of this mechanism, the visual embeddings $\boldsymbol{X}$ pass through a linear layer of size $D/H$, where $H$ denotes the number of heads, forming the Query ($\boldsymbol{Q}_{h}=\{\boldsymbol{q}_{n}\}_{n=1}^{N}$) matrix. The text embeddings $\boldsymbol{Y}$ pass through two different linear layers of size $D/H$, creating the Key ($\boldsymbol{K}_{h}=\{\boldsymbol{k}_{m}\}_{m=1}^{M}$) and Value ($\boldsymbol{V}_{h}=\{\boldsymbol{v}_{m}\}_{m=1}^{M}$) matrices. The Query and Key matrices undergo a matrix multiplication process that is followed by a softmax layer, formulating the values of the attention matrix ($\boldsymbol{A}_{h}=softmax(\boldsymbol{Q}_{h}\times\boldsymbol{K}_{h}^{T})\ and\ \boldsymbol{A}_{h}=\{a_{i,j}\}_{i=1,N}^{j=1,M}$). This matrix goes through a dropout layer and the outcome is multiplied with the Value matrix, defining the output of the $h^{th}$ attention head ($\boldsymbol{Z}_{h}=\{\boldsymbol{z_{n}}\}_{n=1}^{N}\ and\ \boldsymbol{z_{n}}=\{z_{r}\}_{r=1}^{D/H}$). The obtained representations in the output of all the attention heads are then concatenated, and the resulting representations pass through a linear layer of size $D$ and a positional encoding mechanism (similar to the one in [30]), formulating the overall cross-modal embeddings ($\boldsymbol{Z}=\{\boldsymbol{z}_{n}\}_{n=1}^{N}$) in the output of the attention mechanism. Subsequently, as depicted in Fig. 2, these embeddings undergo a dropout and normalization process and then are given as input to a trainable Transformer-based scoring mechanism which computes frame-level importance scores ($\boldsymbol{f}=\{f_{n}\}_{n=1}^{N}$). These scores are finally used by a frame/fragment selection component that forms the video summary based on a predefined temporal fragmentation of the video and a time-budget about the summary duration.
Figure 3: The processing pipeline of the utilized multi-head cross-modal attention mechanism. The visual embeddings form the Query matrix and the text embeddings form the Key and Value matrices based on trainable linear layers. The Query and Key are then multiplied and the computed values pass through a softmax layer, forming the Attention matrix. This matrix is multiplied with the Value, producing the output of an attention head. The outputs of all attention heads are concatenated, and the obtained representations pass through a linear layer and a positional encoding mechanism, formulating the overall cross-modal embeddings in the output of the attention mechanism.
@@
So, in terms of network architecture, SD-VSum is most closely related with the CLIP-It architecture [6] for query-based and multimodal generic video summarization. CLIP-It uses a language-guided multi-head attention mechanism to fuse information from the visual and text modalities, and employs a Transformer-based scorer for assigning importance scores to the video frames. With respect to the representation of the input data, the video is represented using similar (CLIP-based) frame-level embeddings, but the query (in the case of query-based video summarization) or the dense video captions (in the case of multimodal generic video summarization) is represented using a condensed single-vector embedding. The latter is produced by a trainable linear layer that gets as input a vector formed by concatenating a fixed number of sampled (CLIP-based) embeddings of the input text, and projects it into a more condensed representation (of dimension $1\times 512$). As a consequence, the language-guided attention of CLIP-It encodes the dependence between the video frames and the user query in a more compact, and thus less informative, manner. On the contrary, the cross-modal attention mechanism of the proposed SD-VSum network architecture employs a more detailed representation of the input text at the sentence level, allowing SD-VSum to encode the frames’ dependence with different parts of the user script in a more effective way (using an $N\times M$ attention matrix, as shown in Fig. 3, rather than a more condensed $N\times 1$ matrix, as in CLIP-It). This main distinction leads to a noticeably advanced summarization performance, as documented in Section 5.3.
4.2 Training regime
@@
Training of SD-VSum is performed in a supervised learning setting, comparing its output for a given pair of full-length video and user script, with the associated ground-truth summary (i.e., the one used to generate this particular script). In particular, we compute the Binary Cross Entropy (BCE) between the computed scores for the frames of the input video by SD-VSum ($\boldsymbol{f}=\{f_{n}\}_{n=1}^{N}$), and the assigned binary labels to these frames by a human annotator. The computed training loss is then back-propagated to compute the gradients and update all the different trainable parts of the network architecture. The use of the multiple available ground-truth summaries for each video of the dataset as individual training samples, practically increases tenfold the amount of training, validation and test samples, reported in Section 3.2.
@@
A different training regime is followed when evaluating the performance of SD-VSum on the task of multimodal generic video summarization. In this case, the computed frame-level scores for a given pair of full-length video and textual description of its content, are compared with a single ground-truth summary that is formed by averaging the multiple human-annotated summaries for this video, at the frame-level. So, instead of comparing the output of SD-VSum with multiple different ground-truth annotations for the input video using BCE (a contradictory approach that could prevent the training of the network architecture), we calculate a single loss value by using an average ground-truth summary and the Mean Squared Error (MSE).
5 Experiments
5.1 Evaluation protocol
@@
Following the evaluation approach in [9], we generate the video summary for a given input video and user script, by selecting the top-15% scoring frames by a video summarization model. Then, we quantify the similarity between the machine-generated and the ground-truth summary using the F-Score (as percentage). So, a given test video is matched with each one of the multiple available user scripts for it, and each one of the generated summaries is compared with the corresponding ground-truth summary. Through this process, we compute an F-Score for each pair of compared summaries and we average these scores to form the final F-Score for this video. After performing this for all test videos, we calculate the mean of the obtained F-Score values and result in an average score that indicates the method’s performance on the test set.
@@
When evaluating the performance on multimodal generic video summarization we also use the evaluation protocol from [31]. In particular, we measure the alignment between the computed frame-level importance scores for a given video by a video summarization model, and the ground-truth frame-level importance scores for this video (obtained by averaging its multiple available binary ground-truth summaries, at the frame-level), using the Kendall’s $\tau$ [32] and Spearman’s $\rho$ [33] rank correlation coefficients. The computed $\tau$ and $\rho$ values for all test videos are then averaged, defining the performance of the video summarization model on the test set.
@@
The selection of a well-trained model for evaluation on the videos of the test is made based on the recorded performance on the videos of the validation set. More specifically, after each training epoch we measure the performance of the trained model on the videos of the validation set. When then training process is completed, we keep the model with the highest validation-set performance in terms of F-Score, and assess its performance on the test set using the evaluation protocols described above.
5.2 Implementation details

### v2

@@ 3.1. Problem formulation
Let us assume an input video of $N$ frames and a user-provided script (a free-form textual description) of $M$ sentences outlining the content of the desired video summary; different sentences of the script may refer to different parts of the full-length video with varying visual and semantic content. Then, script-driven video summarization aims to select a subset of frames or fragments of the video that are both relevant to one or more sentences of the user-provided script, and important for an overall understanding of the video. The selected frames/fragments should form a video summary that does not exceed a predefined length $L$, which is typically set to $15\%$ of the full-length video duration (Apostolidis et al., 2021a).
@@
Another task that aims to spot parts of the video that are relevant to a freeform query is the task of temporal sentence grounding in videos (Lan et al., 2023) (a.k.a. video grounding). Nevertheless, this task tries to localize one target segment from the full-length video, that corresponds to a given sentence query. On the contrary, script-driven video summarization aims to generate a concise summary, and in this process it needs to select multiple video fragments that relate to the complex query, as well as other less relevant parts that are necessary for providing a synopsis of the full-length video.
3.2. Script-driven video summarization dataset
Figure 2. A keyframe-based representation of the video “v_3OcAjx8e4LU”, the textual annotation in the original VideoXum dataset, and examples of the generated scripts for two ground-truth summaries of this video in the S-VideoXum dataset.
@@ 3.2. Script-driven video summarization dataset
To overcome the observed weaknesses of small video summarization datasets, related to the sensitivity of the evaluation outcomes to different data splits (Apostolidis et al., 2020) and the inability to select a well-trained model in the absence of a validation set (Apostolidis et al., 2021a), we focused on the VideoXum large-scale dataset for cross-modal video summarization (Lin et al., 2024). VideoXum includes $14,000$ open-domain videos up to $12.5$ min. long with diverse visual content, from the ActivityNet Captions dataset (Krishna et al., 2017). Each video is associated with multiple ground-truth video summaries - in the form of frame-level binary scores which denote the inclusion or not of a frame in the video summary - that were obtained by $40$ different human annotators. The dense video captions from ActivityNet Captions, that are considered as text summaries in (Lin et al., 2024), provide a high-level description of the full-length video (see Fig. 2) and thus are more appropriate for multimodal generic video summarization purposes. The existence of multiple ground-truth summaries per video ($10$ in total) makes VideoXum well-suited for the script-driven video summarization task, as it allows to train a summarization method to generate different summaries for a given video, driven by the provided script about the content of each summary. Furthermore, its significantly large size allows to split the available videos into train, validation and test set. Such a data split supports the extensive training of a model using a very wide set of training samples, the selection of a well-trained model that exhibits good generalization capacity through the use of an adequately big validation set, and the robust evaluation of its performance using a large number of testing samples.
@@
To make VideoXum suitable for training and evaluation of script-driven video summarization methods, we extended it by producing natural language descriptions (a.k.a. scripts) of the different available human-annotated ground-truth summaries per video. For this, we employed the publicly-available [^1] https://huggingface.co/llava-hf/LLaVA-NeXT-Video-7B-hf SOTA Large Multimodal Model LLaVA-NeXT-Video-7B (Li et al., 2024). We chose this model because of its strong video understanding capabilities, outperforming other open-source models (e.g., LLaMA-VID) and performing comparably to Gemini Pro [^2] https://llava-vl.github.io/blog/2024-04 -30-llava-next-video/. To align with the adopted video representation strategy in (Lin et al., 2024), we processed each ground-truth summary using one frame per second and obtained the scripts using the following prompt: “Describe the important scenes in the video”. The maximum number of generated tokens was set to $200$, and a $4$-bit quantization was applied to reduce computational cost. The obtained scripts are diverse. Taking into account one summary per video (the first of each video’s ten summaries), the scripts vary across videos from very concise to extensive ones (#words: min = $4$, max = $220$, avg = $126.5$, std = $30.0$; #sentences: min = $1$, max = $32$, avg = $6.1$, std = $1.5$). To estimate the within-video diversity of summaries, we calculated the above statistics separately for each video across its ten summaries, and then averaged over the dataset. The average std of #words and #sentences are $21.7$ and $1.1$, i.e., within-video diversity is also quite high. Visual inspection of sample scripts verified that they are quite diverse and detailed, can refer to multiple unrelated actions, and might contain abstract cues and narrative elements (e.g., see Fig. 2). In addition, the annotations in S-VideoXum correspond to individual ground-truth summaries, rather than the full-length video (as in VideoXum), and focus on different elements of the summary generated by each different annotator, as depicted in Fig. 2. Finally, to allow performance comparisons with multimodal approaches for generic video summarization that use the same data modalities, we produced natural language descriptions of the full-length videos, as well. Once again, we sampled these videos keeping one frame per second and generated descriptions using LLaVA-NeXT-Video-7B and the same prompt.
@@
Since some of the videos from the VideoXum dataset were not publicly-available during this work, the extended S-VideoXum dataset for script-driven video summarization differs from the original one in terms of number of videos, including data for $11,908$ videos. However, a statistical analysis on the obtained videos showed that our dataset maintains the same distribution of train, validation and test samples with VideoXum. In particular, we use $6,782$ samples for training, $1,707$ for validation and $3,419$ for testing.
Prior to constructing the S-VideoXum dataset, we experimented with a small proprietary dataset (S-NewsVSum) of a Media company, including $45$ videos from news broadcast and the associated professionally-edited video summaries and user-authored scripts.
Both datasets are shared at: https://github.com/IDT-ITI/SD-VSum
4. Proposed Approach
4.1. Network architecture
@@
An overview of the SD-VSum network architecture is provided in Fig. 3. Assuming an input video of $N$ frames (following sampling), and a user script of $M$ sentences that outlines the content of the desired summary, SD-VSum produces a video summary conditioned on the script and a time-budget about its length. A set of visual embeddings ($\boldsymbol{X}=\{\boldsymbol{x_{n}}\}_{n=1}^{N}$) of size $D$ ($\boldsymbol{x_{n}}=\{x_{r}\}_{r=1}^{D}$) are obtained for the video frames, using the visual encoder of a pretrained multimodal embedding model. Similarly, a set of text embeddings ($\boldsymbol{Y}=\{\boldsymbol{y_{m}}\}_{m=1}^{M}$) of the same size ($\boldsymbol{y_{m}}=\{y_{r}\}_{r=1}^{D}$) are computed for the different sentences of the user script (one per sentence), using the text encoder of the aforementioned model. These embeddings are then given as input to a multi-head cross-modal attention mechanism. As shown in the upper part of Fig. 3, given the $h^{th}$ attention head of this mechanism, the visual embeddings $\boldsymbol{X}$ pass through a linear layer of size $D/H$, where $H$ denotes the number of heads, forming the Query ($\boldsymbol{Q}_{h}=\{\boldsymbol{q}_{n}\}_{n=1}^{N}$) matrix. The text embeddings $\boldsymbol{Y}$ pass through two different linear layers of size $D/H$, creating the Key ($\boldsymbol{K}_{h}=\{\boldsymbol{k}_{m}\}_{m=1}^{M}$) and Value ($\boldsymbol{V}_{h}=\{\boldsymbol{v}_{m}\}_{m=1}^{M}$) matrices. The Query and Key matrices undergo a matrix multiplication and a softmax process, formulating the values of the attention matrix ($\boldsymbol{A}_{h}=softmax(\boldsymbol{Q}_{h}\times\boldsymbol{K}_{h}^{T})\ and\ \boldsymbol{A}_{h}=\{a_{i,j}\}_{i=1,N}^{j=1,M}$). This matrix goes through a dropout layer and the outcome is multiplied with the Value matrix, defining the output of the $h^{th}$ attention head ($\boldsymbol{Z}_{h}=\{\boldsymbol{z_{n}}\}_{n=1}^{N}\ and\ \boldsymbol{z_{n}}=\{z_{r}\}_{r=1}^{D/H}$). The obtained representations in the output of all attention heads are then concatenated, and pass through a linear layer of size $D$ and a positional encoding mechanism (similar to the one in (Apostolidis et al., 2021b)), forming the overall cross-modal embeddings ($\boldsymbol{Z}=\{\boldsymbol{z}_{n}\}_{n=1}^{N}$) in the output of the attention mechanism. Following, as shown in the lower part of Fig. 3, these embeddings undergo a dropout and normalization process and then are given as input to a trainable Transformer-based scoring mechanism which computes frame-level importance scores ($\boldsymbol{f}=\{f_{n}\}_{n=1}^{N}$). These scores are finally used by a frame/fragment selection component that forms the video summary based on a temporal fragmentation of the video and a time-budget about the summary duration.
Figure 3. Overview of SD-VSum. Given an input video and a user script outlining the content of the desired summary, SD-VSum generates a summary conditioned on the script. Visual and text embeddings are obtained using a pretrained multimodal embedding model, and fused using a trainable cross-modal attention mechanism, as shown in the upper part. The cross-modal embeddings in the output of this mechanism are forwarded to a trainable Transformer-based frame scoring mechanism, which computes frame-level importance scores. These scores are finally used by a frame/fragment selection component that forms the video summary based on a predefined temporal fragmentation of the video and a time-budget about the summary duration.
@@
So, in terms of network architecture, SD-VSum is most closely related with CLIP-It (Narasimhan et al., 2021), which uses a language-guided multi-head attention mechanism to fuse information from the visual and text modalities, and employs a Transformer-based scorer for assigning importance scores to the video frames. With respect to the representation of the input data, the video is represented using similar (CLIP-based) frame-level embeddings, but the input text (query or dense video captions) is represented using a condensed single-vector embedding. The latter is produced by a trainable linear layer that gets as input a vector formed by concatenating a fixed number of sampled (CLIP-based) embeddings of the input text, and projects it into a more condensed representation (of dimension $1\times 512$). As a consequence, the language-guided attention of CLIP-It encodes the dependence between the video frames and the user query in a more compact, and thus less informative, manner. On the contrary, the cross-modal attention mechanism of the proposed SD-VSum network architecture employs a more detailed representation of the input text at the sentence level, allowing SD-VSum to encode the frames’ dependence with different parts of the user script in a more effective way (using an $N\times M$ attention matrix, as shown in the upper part of Fig. 3, rather than a more condensed $N\times 1$ matrix, as in CLIP-It). This main distinction leads to a noticeably advanced summarization performance, as documented in Section 5.3.
4.2. Training regime
@@
Training of SD-VSum is performed in a supervised learning setting, comparing its output for a given pair of full-length video and user script, with the associated ground-truth summary (i.e., the one used to generate this particular script). Specifically, we calculate the Binary Cross Entropy (BCE) loss between the computed scores for the frames of the input video ($\boldsymbol{f}=\{f_{n}\}_{n=1}^{N}$) by SD-VSum, and the assigned binary labels to these frames by an annotator. The calculated loss value is then back-propagated and used to update the trainable parts of the network. The use of the multiple available ground-truth summaries for each video of the dataset as individual training samples, practically increases tenfold the amount of train, validation and test samples, reported in Section 3.2.
@@
A different training regime is adopted when evaluating SD-VSum on multimodal generic video summarization. In this case, the computed frame-level scores for a given pair of full-length video and textual description of its content, are compared with a single ground-truth summary that is formed by averaging the multiple human-annotated summaries for this video, at the frame-level. So, instead of comparing the output of SD-VSum with multiple different ground-truth annotations using BCE (a contradictory approach that could prevent network’s training), we compute a single loss value using an average ground-truth summary and the Mean Squared Error.
5. Experiments
5.1. Evaluation protocol
@@
Following the evaluation approach in (Lin et al., 2024), we generate the video summary for a given input video and user script, by selecting the top-15% scoring frames by a video summarization model. Then, we quantify the similarity between the machine-generated and the ground-truth summary using the F-Score (as percentage). So, a given test video is matched with each one of the multiple available user scripts for it, and each one of the generated summaries is compared with the corresponding ground-truth summary. Through this process, we compute an F-Score for each pair of compared summaries and we average these scores to form the final F-Score for this video. After performing this for all test videos, we calculate the mean of the obtained F-Score values and result in an average score that indicates the model’s performance on the test set.
@@
When evaluating the performance on multimodal generic video summarization we also use the evaluation protocol from (Otani et al., 2019). In particular, we measure the alignment between the computed frame-level importance scores for a given video by a video summarization model, and the ground-truth frame-level importance scores for this video (obtained by averaging its multiple available binary ground-truth summaries, at the frame-level), using the Kendall’s $\tau$ (Kendall, 1945) and Spearman’s $\rho$ (Kokoska and Zwillinger, 2000) rank correlation coefficients. The computed $\tau$ and $\rho$ values for all test videos are then averaged, defining the performance of the video summarization model on the test set.
@@
The selection of a well-trained model for evaluation on the videos of the test is made based on the recorded performance on the videos of the validation set. More specifically, after each training epoch we measure the performance of the trained model on the videos of the validation set. When the training process is completed, we keep the model with the highest validation-set performance in terms of F-Score, and assess its performance on the test set using the evaluation protocols described above.
For the small S-NewsVSum dataset, we follow the same general approach, the difference being that we have only one ground-truth summary per video. Additionally, this being a very small dataset, we perform 5-fold cross validation and report the average performance.
5.2. Implementation details